# A generalization of Barannikov-Kontsevich theorem

### v1

@@ 1.2 A basic question
Is the spectral sequence for $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X}[\![\lambda]\!]_{f}$ degenerates at the $E_{1}$-level?
@@
The spectral sequence degenerates at the $E_{1}$-level if and only if the following equivalent conditions are satisfied.
@@ 1.2.1 Classical results related to the question
There are two classical results related to this question. One appeared in the classical Hodge theory. As Deligne observed, if $X$ is projective and $f=0$, the desired $E_{1}$-degeneration follows from the Hodge decomposition of $H^{\ast}(X,{\mathbb{C}})$. Deligne generalized it to more general algebraic varieties by using mixed Hodge theory. This $E_{1}$-degeneration and its generalizations are not only deep results in the Hodge theory, but also useful for various applications including some vanishing theorems. Deligne and Illusie gave an alternative proof of the $E_{1}$-degeneration using the reduction to the positive characteristic.
@@ 1.3 Barannikov-Kontsevich Theorems and variations
holds for any $j$ and for any complex number $\lambda$.
Theorem 1.5 implies that $H^{j}(X^{\mathop{\rm alg}\nolimits},\Omega_{X^{\mathop{\rm alg}\nolimits}}^{\bullet}[\lambda]_{f})$ are free ${\mathbb{C}}[\lambda]$-modules, and that the $E_{1}$-degeneration of the spectral sequence for the filtered complex $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X^{\mathop{\rm alg}\nolimits}}[\lambda]_{f}$. In the setting of Theorem 1.4, there exists a projective morphism of algebraic varieties $F:Y\to{\mathbb{C}}$ with an open embedding $\iota:X\to Y$ such that $f=F\circ\iota$. Under the assumption that $\mathop{\rm Cr}\nolimits(f)$ is compact, the set $\mathop{\rm Cr}\nolimits(F)$ of critical points of $F$ is decomposed as $\mathop{\rm Cr}\nolimits(F)=\mathop{\rm Cr}\nolimits(f)\sqcup\bigl{(}\mathop{\rm Cr}\nolimits(F)\cap(Y\setminus X)\bigr{)}$. Then, we obtain the $E_{1}$-degeneration for $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X}[\![\lambda]\!]_{f}$ from the $E_{1}$-degeneration for $\mathcal{F}_{\lambda}\Omega^{\bullet}_{Y^{\mathop{\rm alg}\nolimits}}[\lambda]_{f}$.
@@ Acknowledgement
I thank Claude Sabbah for helpful discussions and for his kindness.
@@ 2.1.1 $\mathcal{R}_{X}$-triples
We set ${\boldsymbol{S}}=\bigl{\{}\lambda\in{\mathbb{C}}\,\big{|}\,|\lambda|=1\bigr{\}}$. Let $\sigma:{\boldsymbol{S}}\to{\boldsymbol{S}}$ be defined by $\sigma(\lambda)=-\lambda$. Let $\mathfrak{D}\mathfrak{b}_{{\boldsymbol{S}}\times X/{\boldsymbol{S}}}$ denote the sheaf of distributions on ${\boldsymbol{S}}\times X$ which are continuous with respect to $\boldsymbol{S}$}. (See [16, §0.5].) It is naturally an $\mathcal{R}_{X|{\boldsymbol{S}}\times X}\otimes_{\mathcal{O}_{{\mathbb{C}}_{\lambda}|{\boldsymbol{S}}}}\sigma^{-1}(\mathcal{R}_{X|{\boldsymbol{S}}\times X})$-module by the action $(P_{1}\otimes\sigma^{-1}(P_{2}))\cdot\tau=P_{1}\overline{\sigma^{-1}(P_{2})}\tau$. A sesqui-linear pairing of $\mathcal{R}_{X}$-modules $\mathcal{M}^{\prime}$ and $\mathcal{M}^{\prime\prime}$ is a morphism of $\mathcal{R}_{X|{\boldsymbol{S}}\times X}\otimes_{\mathcal{O}_{{\mathbb{C}}_{\lambda}|{\boldsymbol{S}}}}\sigma^{-1}(\mathcal{R}_{X|{\boldsymbol{S}}\times X})$-modules $\mathcal{M}^{\prime}_{|{\boldsymbol{S}}\times X}\otimes_{\mathcal{O}_{{\mathbb{C}}_{\lambda}|{\boldsymbol{S}}}}\sigma^{-1}(\mathcal{M}^{\prime\prime}_{|{\boldsymbol{S}}\times X})\to\mathfrak{D}\mathfrak{b}_{{\boldsymbol{S}}\times X/{\boldsymbol{S}}}$. Such a tuple $(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C)$ is called an $\mathcal{R}_{X}$-triple.
@@ 2.1.2 Direct image of $\mathcal{R}_{X}$-triples
For any $\mathcal{R}_{X}$-triple $\mathcal{T}=(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C)$, we obtain the $\mathcal{R}_{X}$-triples $F_{\dagger}^{j}(\mathcal{T})=(F_{\dagger}^{-j}(\mathcal{M}^{\prime}),F_{\dagger}^{j}(\mathcal{M}^{\prime\prime}),F_{\dagger}C)$ $(j\in{\mathbb{Z}})$ on $Y$. (See [16, §1.4].) When $\mathcal{T}$ is equipped with an increasing filtration $W$, let $W_{j+\ell}(F_{\dagger}^{j}(\mathcal{T}))$ denote the image of $F_{\dagger}^{j}(W_{\ell}\mathcal{T})\to F_{\dagger}^{j}(\mathcal{T})$ by which we obtain the filtration $W$ on $F_{\dagger}^{j}(\mathcal{T})$.
@@ 2.1.4 Mixed twistor $\mathcal{D}$-modules
A mixed twistor $\mathcal{D}$-module on $X$ is a filtered $\mathcal{R}_{X}$-triple $(\mathcal{T},W)$ and satisfying some additional conditions. (See [11, §7].) Let $\mathop{\rm MTM}\nolimits(X)$ denote the category of mixed twistor $\mathcal{D}$-modules on $X$. The following theorem is fundamental.
@@ 2.2 Integrable mixed twistor $\mathcal{D}$-modules
(See [11, §2.1.5] for integrable $\mathcal{R}_{X}$-triples, which originally goes back to [16].) An integrable $\mathcal{R}_{X}$-triple is called $\widetilde{\mathcal{R}}_{X}$-triple. A morphism of $\widetilde{\mathcal{R}}_{X}$-triples $\mathcal{T}_{i}=(\mathcal{M}^{\prime}_{i},\mathcal{M}^{\prime\prime}_{i},C)$ $(i=1,2)$ is defined to be a morphism $(\varphi^{\prime},\varphi^{\prime\prime})$ of $\mathcal{R}_{X}$-triples such that $\varphi^{\prime}$ and $\varphi^{\prime\prime}$ are $\widetilde{\mathcal{R}}_{X}$-homomorphisms. For a morphism of complex manifolds $F:X\to Y$ and for an $\widetilde{\mathcal{R}}_{X}$-triple $\mathcal{T}$, the $\mathcal{R}_{Y}$-triples $F^{j}_{\dagger}(\mathcal{T})$ are naturally $\widetilde{\mathcal{R}}_{Y}$-triples.
An integrable mixed twistor $\mathcal{D}$-module on $X$ is a filtered $\widetilde{\mathcal{R}}_{X}$-triple $(\mathcal{T},W)$ satisfying some conditions. (See [11, §7.2.3].) Let $\mathop{\rm MTM}\nolimits_{\mathop{\rm reg}\nolimits}^{\mathop{\rm int}\nolimits}(X)$ denote the category of integrable mixed twistor $\mathcal{D}_{X}$-modules whose underlying mixed twistor $\mathcal{D}_{X}$-modules are regular. Let $\mathcal{C}_{\mathop{\rm reg}\nolimits}(X)$ denote the full subcategory of $\widetilde{\mathcal{R}}_{X}$-modules underlying regular integrable mixed twistor $\mathcal{D}_{X}$-modules, i.e., an $\widetilde{\mathcal{R}}_{X}$-module $\mathcal{M}^{\prime\prime}$ is an object of $\mathcal{C}_{\mathop{\rm reg}\nolimits}(X)$ if and only if there exists $((\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C),W)\in\mathop{\rm MTM}\nolimits^{\mathop{\rm int}\nolimits}_{\mathop{\rm reg}\nolimits}(X)$.
@@ 2.3 $\widetilde{\mathcal{R}}_{X}$-modules induced by Hodge modules
Proof  There exists a natural functor from the category of mixed Hodge modules on $X$ to $\mathop{\rm MTM}\nolimits^{\mathop{\rm int}\nolimits}_{\mathop{\rm reg}\nolimits}(X)$ as explained in [11, §13.5]. In the level of filtered $\mathcal{D}$-modules, it is given as above.
@@ 3.1.3 $E_{1}$-degeneration and long exact sequences
We recall that the $E_{1}$-degeneration condition of a spectral sequence associated with a filtered complex for the convenience of readers. Let $\mathcal{A}$ be an abelian category. Let $(K^{\bullet},d)$ be a complex with a decreasing filtration $F^{\bullet}(K^{\bullet})$. To simplify the notation, we set $F^{p,p+r}(K^{j}):=F^{p}(K^{j})/F^{p+r}(K^{j})$.
@@
$H^{j}(F^{p,p+r}(K^{\bullet}))\to H^{j}(F^{p,p+1}(K^{\bullet}))$ are epimorphisms for any $j,p\in{\mathbb{Z}}$ and $0\leq r\leq r_{0}$.
@@
$H^{j}(F^{p,p+r}(K^{\bullet}))\to H^{j}(F^{p,p+r-1}(K^{\bullet}))$ are epimorphisms for any $j,p\in{\mathbb{Z}}$ and $0\leq r\leq r_{0}$.
Proof  It is easy to see that $B(r_{0})$ implies $A(r_{0})$. Suppose that $A(r_{0})$ holds. We shall prove $H^{j}(F^{p,p+r}(K^{\bullet}))\to H^{j}(F^{p,p+r-1}(K^{\bullet}))$ are epimorphisms for any $j,p\in{\mathbb{Z}}$ and $0\leq r\leq r_{0}$, by an induction on $r$. We have the following commutative diagram:
@@
By $A(r_{0}-1)$, $c$ is a monomorphism, and we have $b\circ a=0$. There exists a unique morphism $\varphi:H^{j}(F^{p,p+1})\to H^{j+1}(F^{p+r_{0}-1,p+r_{0}})$ such that $c\circ\varphi=b$. By the construction, $\varphi=d_{r_{0}-1}$. Hence, $A(r_{0})$ holds if and only if $d_{r_{0}-1}=0$.
@@
The $E_{1}$-degeneration holds if and only if one of the following equivalent conditions holds.
@@ 3.2.1 Main Theorem
Let us explain a refined statement of Theorem 1.6. We define the filtration $\mathcal{F}_{\lambda}^{k}(\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})=\lambda^{k}(\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})$ for $k\in{\mathbb{Z}}_{\geq 0}$. Note that $\mathop{\rm Gr}\nolimits_{\mathcal{F}_{\lambda}}^{k}\bigl{(}\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\bigr{)}$ is isomorphic to $\iota_{\lambda}^{\ast}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})$ for any $k\geq 0$. Let us state the main theorem of this paper.
@@
The $E_{1}$-degeneration holds for the filtered complex $\mathcal{F}_{\lambda}(\mathcal{M}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})$ with respect to the push-forward by the projection $\mathcal{X}\to{\mathbb{C}}_{\lambda}$. In other words, we obtain the following exact sequences for any $\ell$ and $j$:

### v2

@@ 1.2 A basic question
Does the spectral sequence for $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X}[\![\lambda]\!]_{f}$ degenerate at the $E_{1}$-level?
@@
The spectral sequence degenerates at the $E_{1}$-level if and only if one of the following equivalent conditions are satisfied.
@@ 1.2.1 Classical results related to the question
There are two classical results related to this question. One appeared in the classical Hodge theory. As Deligne observed, if $X$ is projective and $f=0$, the desired $E_{1}$-degeneration follows from the Hodge decomposition of $H^{\ast}(X,{\mathbb{C}})$. Deligne generalized it to more general algebraic varieties by using mixed Hodge theory. This $E_{1}$-degeneration and its generalizations are not only deep results in the Hodge theory, but also useful for various applications including some vanishing theorems. Deligne and Illusie [2] gave an alternative proof of the $E_{1}$-degeneration using the reduction to the positive characteristic.
@@ 1.3 Barannikov-Kontsevich Theorems and variations
holds for any $j$.
Theorem 1.5 implies that $H^{j}(X^{\mathop{\rm alg}\nolimits},\Omega_{X^{\mathop{\rm alg}\nolimits}}^{\bullet}[\lambda]_{f})$ are free ${\mathbb{C}}[\lambda]$-modules, and the $E_{1}$-degeneration of the spectral sequence for the filtered complex $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X^{\mathop{\rm alg}\nolimits}}[\lambda]_{f}$. In the setting of Theorem 1.4, there exists a projective morphism of algebraic varieties $F:Y\to{\mathbb{C}}$ with an open embedding $\iota:X\to Y$ such that $f=F\circ\iota$. Under the assumption that $\mathop{\rm Cr}\nolimits(f)$ is compact, the set $\mathop{\rm Cr}\nolimits(F)$ of critical points of $F$ is decomposed as $\mathop{\rm Cr}\nolimits(F)=\mathop{\rm Cr}\nolimits(f)\sqcup\bigl{(}\mathop{\rm Cr}\nolimits(F)\cap(Y\setminus X)\bigr{)}$. Then, we obtain the $E_{1}$-degeneration for $\mathcal{F}_{\lambda}\Omega^{\bullet}_{X}[\![\lambda]\!]_{f}$ from the $E_{1}$-degeneration for $\mathcal{F}_{\lambda}\Omega^{\bullet}_{Y^{\mathop{\rm alg}\nolimits}}[\lambda]_{f}$.
@@ Acknowledgement
I thank Claude Sabbah for helpful discussions and for his kindness. I also thank him for suggesting the improvement of this manuscript.
@@ 2.1.1 $\mathcal{R}_{X}$-triples
We set ${\boldsymbol{S}}=\bigl{\{}\lambda\in{\mathbb{C}}\,\big{|}\,|\lambda|=1\bigr{\}}$. Let $\sigma:{\boldsymbol{S}}\to{\boldsymbol{S}}$ be defined by $\sigma(\lambda)=-\lambda$. Let $\mathfrak{D}\mathfrak{b}_{{\boldsymbol{S}}\times X/{\boldsymbol{S}}}$ denote the sheaf of distributions on ${\boldsymbol{S}}\times X$ which are continuous with respect to $\boldsymbol{S}$}. (See [17, §0.5].) It is naturally an $\mathcal{R}_{X|{\boldsymbol{S}}\times X}\otimes_{\mathcal{O}_{{\mathbb{C}}_{\lambda}|{\boldsymbol{S}}}}\sigma^{-1}(\mathcal{R}_{X|{\boldsymbol{S}}\times X})$-module by the action $(P_{1}\otimes\sigma^{-1}(P_{2}))\cdot\tau=P_{1}\overline{\sigma^{-1}(P_{2})}\tau$. A sesqui-linear pairing of $\mathcal{R}_{X}$-modules $\mathcal{M}^{\prime}$ and $\mathcal{M}^{\prime\prime}$ is a morphism of $\mathcal{R}_{X|{\boldsymbol{S}}\times X}\otimes_{\mathcal{O}_{{\mathbb{C}}_{\lambda}|{\boldsymbol{S}}}}\sigma^{-1}(\mathcal{R}_{X|{\boldsymbol{S}}\times X})$-modules $\mathcal{M}^{\prime}_{|{\boldsymbol{S}}\times X}\otimes_{\mathcal{O}_{{\mathbb{C}}_{\lambda}|{\boldsymbol{S}}}}\sigma^{-1}(\mathcal{M}^{\prime\prime}_{|{\boldsymbol{S}}\times X})\to\mathfrak{D}\mathfrak{b}_{{\boldsymbol{S}}\times X/{\boldsymbol{S}}}$. Such a tuple $(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C)$ is called an $\mathcal{R}_{X}$-triple.
@@ 2.1.2 Direct image of $\mathcal{R}_{X}$-triples
For any $\mathcal{R}_{X}$-triple $\mathcal{T}=(\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C)$, we obtain the $\mathcal{R}_{X}$-triples $F_{\dagger}^{j}(\mathcal{T})=(F_{\dagger}^{-j}(\mathcal{M}^{\prime}),F_{\dagger}^{j}(\mathcal{M}^{\prime\prime}),F_{\dagger}C)$ $(j\in{\mathbb{Z}})$ on $Y$. (See [17, §1.4].) When $\mathcal{T}$ is equipped with an increasing filtration $W$, let $W_{j+\ell}(F_{\dagger}^{j}(\mathcal{T}))$ denote the image of $F_{\dagger}^{j}(W_{\ell}\mathcal{T})\to F_{\dagger}^{j}(\mathcal{T})$ by which we obtain the filtration $W$ on $F_{\dagger}^{j}(\mathcal{T})$.
@@ 2.1.4 Mixed twistor $\mathcal{D}$-modules
A mixed twistor $\mathcal{D}$-module on $X$ is a filtered $\mathcal{R}_{X}$-triple $(\mathcal{T},W)$ and satisfying some additional conditions. (See [12, §7].) Let $\mathop{\rm MTM}\nolimits(X)$ denote the category of mixed twistor $\mathcal{D}$-modules on $X$. The following theorem is fundamental.
@@ 2.2 Integrable mixed twistor $\mathcal{D}$-modules
(See [12, §2.1.5] for integrable $\mathcal{R}_{X}$-triples, which originally goes back to [17].) An integrable $\mathcal{R}_{X}$-triple is called $\widetilde{\mathcal{R}}_{X}$-triple. A morphism of $\widetilde{\mathcal{R}}_{X}$-triples $\mathcal{T}_{i}=(\mathcal{M}^{\prime}_{i},\mathcal{M}^{\prime\prime}_{i},C)$ $(i=1,2)$ is defined to be a morphism $(\varphi^{\prime},\varphi^{\prime\prime})$ of $\mathcal{R}_{X}$-triples such that $\varphi^{\prime}$ and $\varphi^{\prime\prime}$ are $\widetilde{\mathcal{R}}_{X}$-homomorphisms. For a morphism of complex manifolds $F:X\to Y$ and for an $\widetilde{\mathcal{R}}_{X}$-triple $\mathcal{T}$, the $\mathcal{R}_{Y}$-triples $F^{j}_{\dagger}(\mathcal{T})$ are naturally $\widetilde{\mathcal{R}}_{Y}$-triples.
An integrable mixed twistor $\mathcal{D}$-module on $X$ is a filtered $\widetilde{\mathcal{R}}_{X}$-triple $(\mathcal{T},W)$ satisfying some conditions. (See [12, §7.2.3].) Let $\mathop{\rm MTM}\nolimits_{\mathop{\rm reg}\nolimits}^{\mathop{\rm int}\nolimits}(X)$ denote the category of integrable mixed twistor $\mathcal{D}_{X}$-modules whose underlying mixed twistor $\mathcal{D}_{X}$-modules are regular. Let $\mathcal{C}_{\mathop{\rm reg}\nolimits}(X)$ denote the full subcategory of $\widetilde{\mathcal{R}}_{X}$-modules underlying regular integrable mixed twistor $\mathcal{D}_{X}$-modules, i.e., an $\widetilde{\mathcal{R}}_{X}$-module $\mathcal{M}^{\prime\prime}$ is an object of $\mathcal{C}_{\mathop{\rm reg}\nolimits}(X)$ if and only if there exists $((\mathcal{M}^{\prime},\mathcal{M}^{\prime\prime},C),W)\in\mathop{\rm MTM}\nolimits^{\mathop{\rm int}\nolimits}_{\mathop{\rm reg}\nolimits}(X)$.
@@ 2.3 $\widetilde{\mathcal{R}}_{X}$-modules induced by Hodge modules
Proof  There exists a natural functor from the category of mixed Hodge modules on $X$ to $\mathop{\rm MTM}\nolimits^{\mathop{\rm int}\nolimits}_{\mathop{\rm reg}\nolimits}(X)$ as explained in [12, §13.5]. In the level of filtered $\mathcal{D}_{X}$-modules and $\widetilde{\mathcal{R}}_{X}$-modules, it is given as above.
@@ 3.1.3 $E_{1}$-degeneration and long exact sequences
We recall that the $E_{1}$-degeneration condition of a spectral sequence associated with a filtered complex for the convenience of readers. Let $(K^{\bullet},d)$ be a complex with a decreasing filtration $F^{\bullet}(K^{\bullet})$ in an abelian category $\mathcal{A}$. To simplify the notation, we set $F^{p,p+r}(K^{j}):=F^{p}(K^{j})/F^{p+r}(K^{j})$.
@@
$H^{j}(F^{p,p+r}(K^{\bullet}))\to H^{j}(F^{p,p+1}(K^{\bullet}))$ are epimorphisms for any $j,p\in{\mathbb{Z}}$ and $1\leq r\leq r_{0}$.
@@
$H^{j}(F^{p,p+r}(K^{\bullet}))\to H^{j}(F^{p,p+r-1}(K^{\bullet}))$ are epimorphisms for any $j,p\in{\mathbb{Z}}$ and $1\leq r\leq r_{0}$.
Proof  It is easy to see that $B(r_{0})$ implies $A(r_{0})$. Suppose that $A(r_{0})$ holds. We shall prove $H^{j}(F^{p,p+r}(K^{\bullet}))\to H^{j}(F^{p,p+r-1}(K^{\bullet}))$ are epimorphisms for any $j,p\in{\mathbb{Z}}$ and $1\leq r\leq r_{0}$, by an induction on $r$. We have the following commutative diagram:
@@
By $A(r_{0}-1)$, $c$ is a monomorphism, and we have $b\circ a=0$. There exists a unique morphism $\varphi:H^{j}(F^{p,p+1})\to H^{j+1}(F^{p+r_{0}-1,p+r_{0}})$ such that $c\circ\varphi=a$. By the construction, $\varphi=d_{r_{0}-1}$. Hence, $A(r_{0})$ holds if and only if $d_{r_{0}-1}=0$.
@@
The $E_{1}$-degeneration holds if and only if the following equivalent condition holds.
@@ 3.2.1 Main Theorem
Let us explain a refined statement of Theorem 1.6. We define the filtration $\mathcal{F}_{\lambda}^{k}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})=\lambda^{k}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})$ for $k\in{\mathbb{Z}}_{\geq 0}$. Note that $\mathop{\rm Gr}\nolimits_{\mathcal{F}_{\lambda}}^{k}\bigl{(}\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}}\bigr{)}$ is isomorphic to $\iota_{\lambda}^{\ast}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})$ for any $k\geq 0$. Let us state the main theorem of this paper.
@@
The $E_{1}$-degeneration holds for the filtered complex $\mathcal{F}_{\lambda}(\mathcal{M}_{f}\otimes\widetilde{\Omega}^{\bullet}_{\mathcal{X}/{\mathbb{C}}})$ with respect to the push-forward by the projection $\mathcal{X}\to{\mathbb{C}}_{\lambda}$. In other words, we obtain the following exact sequences for any $\ell$ and $j$: